### v2

# Rethinking Gating Mechanism in Sparse MoE: Handling Arbitrary Modality Inputs with Confidence-Guided Gate

Abstract
Effectively managing missing modalities is a fundamental challenge in real-world multimodal learning scenarios, where data incompleteness often results from systematic collection errors or sensor failures. Sparse Mixture-of-Experts (SMoE) architectures have the potential to naturally handle multimodal data, with individual experts specializing in different modalities. However, existing SMoE approach often lacks proper ability to handle missing modality, leading to performance degradation and poor generalization in real-world applications. We propose ConfSMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture by taking the opinion of experts and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence. Inspired by our theoretical analysis, ConfSMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth signal. This naturally relieves expert collapse without introducing additional load balance loss function. We show that the insights of expert collapse aligns with other gating mechanism such as Gaussian and Laplacian gate. The proposed method is evaluated on four different real world dataset with three distinct experiment settings to conduct comprehensive analysis of ConfSMoE on resistance to missing modality and the impacts of proposed gating mechanism. We release our official implementation in: https://github.com/IcurasLW/Official-Repository-of-ConfSMoE.git
1 Introduction
As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision supportZheng et al. (2024a, c), cross-modal inferenceWu et al. (2017); Cheng et al. (2024), temporal reasoningZheng et al. (2024b); Ni et al. (2024), and natural language understandingFedus et al. (2022); Masoudnia and Ebrahimpour (2014). Sparse Mixture-of-Experts (SMoE) architecturesYao et al. (2024); Yun et al. (2024); Mustafa et al. (2022) are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation. They are particularly effective when all input modalities are present, enabling diverse feature extraction and flexible expert routing. However, this assumption is often violated in practice. Incomplete modality observations are common in real-world scenarios due to sensor failures, privacy restrictions, or heterogeneous data collection pipelines. When one or more modalities are missing, SMoE models receive partial input signals, which undermines the reliability of the gating mechanism and results in suboptimal expert selection. Although some methods attempt to assign different modality combinations to different experts, this strategy becomes computationally intractable as the number of modalities grows exponentially. To reduce this complexity, recent SMoE frameworks typically activate the Top-K experts per sample. However, the use of softmax-based routingFedus et al. (2022) tends to produce sharp distributions, often concentrating the routing on a single expert. This leads to the expert collapse problem Fedus et al. (2022); Han et al. (2024); Yun et al. (2024), where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization. Efforts to mitigate this issue using load-balancing losses introduce additional optimization challenges. In particular, they may cause unstable or ambiguous expert selection during training, as illustrated in Figure3(a).
(a) Expert Collapse
(b) Imputation by shared
(c) Ours
Figure 1: Expert Collapse of MoE and Comparison with Two Different Modality Imputation
While recent worksFedus et al. (2022); Han et al. (2024); Yun et al. (2024) has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge. In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing Ma et al. (2021); Wang et al. (2023); Yao et al. (2024) as shown in Figure 1(b). However, these methods typically introduce multiple auxiliary losses to jointly learn task-relevant, shared, and modality-specific features, which increases training complexity and may inadvertently erode modality-unique information. To alleviate this issue, recent work has proposed using learnable modality banksYun et al. (2024) to impute missing inputs based on the available modality combinations. Nevertheless, such banks are usually initialized randomly, similar to prompt tuningJia et al. (2022); Lester et al. (2021), and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (ConfSMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities. Our method stems from a core insight: modality-specific structure should be preserved and leveraged for both reconstruction and routing, a property often overlooked in prior work. Based on this, we propose a two-stage imputation strategy that first estimates missing modalities using intra-modality correlations, and then refines these estimates through instance-level cross-modal alignment, as illustrated in Figure 1(c). The rationale behind this approach is inspired by the modality gap phenomenon discussed by Liang et al. (2022), which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks. Thus, it is natural to first impute missing modalities using intra-modality information subsequently enhance them through instance-level cross-modal refinement.
We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalizationFedus et al. (2022); Wang et al. (2024); Yun et al. (2024); Han et al. (2024); Chi et al. (2022). Through theoretical analysis, we show that conventional load-balancing losses exacerbate this issue by introducing gradient conflicts during optimization, reflecting by the "Sinoual Wave" pattern in expert selection plot shown in Figure 3(a). To resolve this, we design a confidence-guided expert routing mechanism that replaces softmax gating with token-level confidence scores. This formulation not only avoids the adverse effects of entropy-based balancing but also enables stable, interpretable expert selection. By decoupling routing from the optimization bottlenecks of prior designs, ConfSMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios. Our contributions are the following:
  • A confidence-guided expert routing mechanism is proposed to mitigate expert collapse by decoupling gating from softmax-induced sharpness. Token-level confidence is used as a interpretable signal for expert selection.
  • A two-stage imputation framework is developed, where missing modalities are first reconstructed using modality-specific patterns, then refined with instance-level cross-modal context to preserve structural fidelity.
  • Empirical results across three missing-modality settings demonstrate strong robustness and generalization. Performance improvements are further supported by ablation studies that align with theoretical analysis.
2 Preliminaries
2.1 Mixture-of-Expert Architecture
In this work, we build upon the SMoE framework by incorporating expert routing into the Transformer’s feedforward layers, enabling dynamic token-to-expert assignment based on learned routing scores. This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting Han et al. (2024), allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction. We begin with the standard Softmax SMoEs with $N$ number of expert as a baseline configuration. Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$ from expert pool $\mathbf{E}=[E_{1},E_{2},\dots,E_{N}]\in\mathbb{R}^{d\times d}$. Then, Top-K function is applied to select the Top-K experts in forward, the final tokens/samples representation is weighted summation from Top-K experts with corresponding routing score $g_{i}$. Formally, we can represent the forward process of SMoE as:
$E(\mathbf{h})=\mathbf{h}+\sum_{i}^{K}g_{i}E_{i}(\mathbf{h})$ (1)
2.2 What Causes Expert Collapse?
The Jacobian matrix of Eq. 1 w.r.t $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
$\mathbf{J}_{MoE}=\sum_{i}^{K}\frac{\partial g_{i}}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
[…]
$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{diag(g)}_{K}-\textbf{g}\textbf{g}_{K}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
2.3 Gradient of Load Balance Loss
This expert collapse phenomenon restrict the model learning capacity since many other expert are wasted and cannot reflect the actual expert diversity. One widely-used solutions to achieve balance expert usage is to introduce additional load balance loss Yun et al. (2024); Han et al. (2024). However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to routing score terms in Eq 3. To show this, we consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a diverse softmax routing distribution, which encourage higher entropy on the routing score distribution. From Assumption 1 defined in Appendix, we define that the objective of any $\mathcal{L}_{load}$ as $\frac{1}{\mathcal{H(\mathbf{g})}}$ over the softmax distribution, where $\mathcal{H(\mathbf{g})}=-\sum_{i=1}^{N}g_{i}log\ g_{i}$ represent the entropy of softmax distribution g. Taking the Jacobian over the $\mathcal{L}_{load}$ will give us (See detailed derivation in Appendix. B.2):
$\mathbf{J}_{load}=[\frac{1}{\mathcal{H}(\textbf{g})^{2}}(log(\textbf{g})+1)^{\top}]\cdot(\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top})$ (4)
Recall that $\mathbf{E}_{K}(\mathbf{h})$ in Eq. 3 is a set of expert with ReLU activation, we assume this network should not output negative value as activation is located on positive side. We also know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always Positive Semi-Definite (PSD) Gao and Pavel (2017), making the optimization of better routing score is always positive. In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{N})+1]$. When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$, are closed to 0. Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value, producing conflict gradient between $\mathbf{J}_{load}$ and the first term of $\mathbf{J}_{MoE}$. The Top-K operation does not change PSD property of term, but the gradient will be backward through the off Top-K term in $J_{MoE}$. It is also worthy to notice that gradient in Eq. 4 conflict occurs for all Top-K number as the load balance loss consider all $g_{i}$ regardless of number K expert selection. Taking an extreme example, Top-1 selection Fedus et al. (2022) would generate one $g_{i}$ out of g for the second term of Eq. 3 while Eq. 4 still take all $g_{i}$ to gradient computation. Therefore, the opposite gradient from axillary load balance loss makes the training difficult to converge to global minimum and Expert selection becomes ambiguous, failing to reflect actual capacity of expert.
3 Propose Method
3.1 ConfNet:Confidence-Guided Gating Network
From Eq. 4 and Eq. 3, we understand that load balance loss shows opposite gradient direction to softmax router in typical SMoE. To avoid this gradient conflict and balance the load, the potential solution is to remove the association of sharp distribution from Eq. 3 and attempt to explicitly achieve load balance. One is to replace the routing function as illustrated in Figure. 2 (c) and one is to decouple the gating score with softmax router as illustrated in Figure. 2 (d). A typical example from a recent work, FuseMoE Han et al. (2024), they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding by $\mathcal{L}_{1}$ distance. Then the Top-K experts are selected with highest similarity to expert embedding and Laplacian router is not a sharp distribution so that the load is explicit balanced. The second approach is to decouple the connection of gating score and expert selection. One simple approach is to consider all expert are equally important, then the final representation is averaged from all selected expert. However, this averaged approach does not emphasize the difference of expert and lose the specification of expert.
Therefore, we propose a Confidence-Guided Gating Network Pool, which introduces a novel routing strategy guided by task confidence. Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity as in Han et al. (2024); Fedus et al. (2022). For each expert $E_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, we introduce an auxiliary confidence estimation network: a one-layer linear network $U_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}^{1}$. $v_{i}$ is then mapped to a confidence score $c_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$. To learn the true confidence $p_{t}$ of downstream task w.r.t $y_{t}$, where $t$ is the notation for ground truth. We apply MSE loss as an axillary loss that minimize the difference between $c_{i}$ and $p_{t}$. Formally, given any classification dataset $D$ with supervised signal $y_{i}$ and number of instance $|D|$, we consider the task loss $\mathcal{L}_{task}$ as Cross Entropy Loss and the final objective is to optimize:
$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}y_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(c_{i}-p_{t})^{2}$ (5)
Note that Eq. 5 is convex as both cross-entropy and MSE loss are well-known convex. During both training and inference, $c_{i}$ serves as a proxy gating score for each expert to token fusion in Eq. 2. The $c_{i}$ is Token-level confidence, which each token has one corresponding $c_{i}$ for aggregation as in $\sum^{K}_{i}G(\mathbf{h})E_{i}(\mathbf{h})$ as in Eq. 1, resulting in a variant of proposed method for fine-grained feature learning, denoted as ConfSMoE-T. Additionally, a expert-level gating signal is also implemented for coarse-grained information retrieval. We aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment. The multimodal interaction is learned during the token assignment for all experts, which all modalities are shared one identical expert pool. This confidence-driven routing enables both fine-grained specialization and robust multimodal interaction, leading to improved generalization and interpretability.
Figure 2: The Framework of ConfSMoE: (a) ConfSMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation. The ConfNet pool is a set of one-layer linear weight $\mathbf{U}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$. Missing modality imputation is achieved by taking the mean of random instances from the corresponding modality of training data as common modality specific feature. The instance-specific feature are sparsely aggregated from existing modality. (b) Sharp gating score causes expert collapse. (c)(d) "Fair" gating score distribution and decoupling gating score can alleviate the expert collapse without the helps of load balance loss.
3.2 Missing Modality Modeling: Two-Stage Imputation
Inspired by the modality gap phenomenon observed in contrastive learning Liang et al. (2022), we hypothesize that a similar gap exists in our multimodal setting, and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4. The presence of this modality gap suggests that imputing a missing modality using only shared information from the available modalities is inherently challenging, due to the distinct representational spaces each modality occupies. Ideally, the imputed modality should preserve the characteristics of its original distribution without being significantly biased by other modalities. To address this, we propose a Two-stage Imputation strategy to reconstruct the representative feature from the missing modality distribution itself, and further refine the imputed modality by adding up instance-specific feature from the available modalities.
Pre-Imputation: To retrieve modality-specific feature, given a sample $x_{i}$ with missing modality $M_{m}\in\mathbb{R}^{s\times d}$ and arbitrary available modality $M_{a}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length. Pre-imputation aims to obtain the modality-specific feature from other instances of $M_{m}$ distribution. One typical way is to take the mean feature distribution of $M_{m}$, but simply taking the mean may be overfitting to training set and all representation across different modality samples will be deterministic and uninformative. In contrast to mean representation, we wish to introduce some stochasticity during the pre-imputation as this stochasticity may prevent the representation to collapse to the same representation. Specifically, we generate an pre-imputed modality embedding $\bar{M}_{m}$ by sampling $n$ number of instances out of the modality pool $M_{m,i}=[M_{m,1},M_{m,2},\cdots,M_{m,n}]$.
$\bar{M}_{m}=\frac{1}{n}\sum^{n}_{i=1}\mathbf{1}M_{m,i}$ (6)
Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{m}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected. $\bar{M}_{m}$ will be sent to ConfSMoE for learning the interaction across different modalities. We denoted $M_{m}^{*}$ as the pre-imputed modality representation after ConfSMoE and $M^{k}_{a}$ as the available modality representation after k-th expert.
Post-Imputation: To refine $M^{*}_{m}$ by available modality, we proposed a post-modality imputation to further capture the interactions between available and missing modality. Unlike conventional approaches that solely rely on the shared representation from one network, our method leverages information from Top-K experts networks by selectively incorporating tokens. This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert to available modality. We employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across $M^{k}_{a}$. By enforcing sparsity, the model is encouraged to attend to the most relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation. Recall from Eq. 1, Top-K MoE generate $K$ specialized representation from available modality $M_{a}^{k}\in\mathbb{R}^{s\times d}$, $k=0,1,\dots K$. Then, we concatenate all $M_{a}^{k}$ returned from $K$ expert to form opinion representation from different expert, denoted as $M_{a}^{*}\in\mathbb{R}^{(s\times K)\times d}$. We apply a sparse cross attention over $M_{m}^{*}$ and $M_{a}^{*}$ to select most relevant tokens to refine $M_{m}^{*}\in\mathbb{R}^{s\times d}$. The SparseCrossAttention takes $M_{m}^{*}$ as Query modality $\mathbf{Q}=\mathbf{W}_{q}M_{m}^{*}$, $M_{a}^{*}$ as Key modality $\mathbf{K}=\mathbf{W}_{k}M_{a}^{*}$ and Value modality $\mathbf{V}=\mathbf{W}_{v}M_{a}^{*}$, where $\mathbf{Q}\in\mathbb{R}^{s\times d}$ and $\mathbf{K},\mathbf{V}\in\mathbb{R}^{(s\times K)\times d}$. Note that $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are shared between all modalities. Then, we obtained the cross attention score $\mathbf{A}\in\mathbb{R}^{s\times(s\times K)}$ from the dot product of $\mathbf{Q}$ and $\mathbf{K}$. Then, we apply a binary mask $\mathbf{M}$ over the attention scores $\mathbf{A}$ to retain only the Top-T entry based on the rank of attention score (1 indicate presence) alongside the dimension of $\mathbf{K}$, yielding a sparse attention map $\mathbf{A}^{*}$. $T=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
$\mathbf{A}^{*}=\mathbf{M}\odot softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}),\in\mathbb{R}^{s\times(s\times K)}$ (7)
Thus, we aggregate the value matrix $V$ using the sparse attention weights $A^{*}$ to obtain the output of $SparseCrossAttention$ (SCA). When multiple available modality $M_{a}^{*}$ are presented in the dataset, we compute their contributions individually and sum them to refine $M_{m}^{*}$:
$M^{*}_{m}=LayerNorm\left(M_{m}^{*}+\sum^{a}SCA(M_{m}^{*},M_{a}^{*},M_{a}^{*})\right)=LayerNorm\left(M_{m}^{*}+\sum^{a}\mathbf{A}^{*}\mathbf{V}\right)$ (8)
In addition, we expect this refinement to redistribute the representation of $M_{m}^{*}$ instead of generating large scale representation and lead to diverge training. A LayerNorm layer is applied to scale the feature map back to consistent distribution as other layers. With SCA, the Post-imputation serves as a refinement module to add up instance-specific feature on $M_{m}^{*}$, which each modality can contribute differently on different sub-segment representation of $M_{m}^{*}$. We visualize the attention weights after imputation in the Appendix. Figures 13, 14, and 15 illustrate these weights, showing that the query modality can dynamically adjust the significance of information by leveraging both expert specialization and shared cross-modal cues. The attention maps reveal distinct weight distributions across different modalities and experts.
4 Experiment
4.1 Experiment Setting
To ensure a fair comparison with baseline methods, we adopt the best-performing hyperparameter settings reported in their respective papers and official implementations. Additionally, we standardize the hidden dimension to 128 across all models. To eliminate confounding effects introduced by varying encoder architectures, we use the same modality-specific encoders for all models, as differences in encoder design can significantly influence final performance. For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT Alsentzer et al. (2019) for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet Cohen et al. (2022) from TorchXRayVision for chest X-ray (CXR) images. Experiment Setting I: Natural Missing Modality. This setting reflects real-world missingness patterns observed in clinical datasets. Experiment Setting II: Random Modality Dropout. A fixed percentage of modalities are randomly dropped for each instance, while ensuring that each instance retains at least one modality in both train and test set. Experiment Setting III: Asymmetric Modality Dropout. Half of the modalities are randomly dropped during training, while testing is performed with only one or two modalities present per instance, simulating domain shifts or deployment constraints. All experiments are 3-Fold cross-validate in different seeds for fair comparison.
4.2 Primary Results
Experiment Setting I: In Table 1 and Appendix Table 8, we provide the experiments results on experiment setting I for three different ICU benchmarking task, where the modalities are natural missing instead of random. We obtain the following observations from the results: (1) In clinical benchmarking tasks 48-IHM, LOS, and 25-PHE, our proposed methods can significantly outperform the most recent baseline approaches such as FuseMoE and FlexMoE and other tradidtional multimodal learning methods in terms of F1 and AUC score. For MIMIC-III, the improvement on F1 and AUC is ranging from 1.81% to 3.91% and 0.92% to 1.22% respectively. (2) The improvement is even further when the size of dataset scale to bigger dataset but the same domain like MIMIC-IV. The improvement on F1 and AUC is ranging from 1.37% to 4.12% and 1.96% to 4.85%. The above empirical evidence roots and supports our claims that the model is strongly robust to missing modality and learn more comprehensive understanding on multimodal inputs.
Table 1: Experiment setting I: Main results on MIMIC-IV
Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
48-IHM F1 39.58 $\pm$ 1.12 32.97 $\pm$ 1.08 45.39 $\pm$ 1.60 11.60 $\pm$ 0.54 43.76 $\pm$ 0.44 30.14 $\pm$ 1.15 40.21 $\pm$ 1.29 35.29 $\pm$ 0.30 49.18 $\pm$ 0.22 48.32 $\pm$ 0.30
AUC 76.60 $\pm$ 0.98 78.29 $\pm$ 0.25 80.43 $\pm$ 0.76 67.08 $\pm$ 0.77 82.97 $\pm$ 0.61 71.34 $\pm$ 0.88 78.05 $\pm$ 0.72 80.45 $\pm$ 0.44 85.24 $\pm$ 0.10 85.09 $\pm$ 0.17
LOS F1 58.52 $\pm$ 0.44 56.22 $\pm$ 0.28 57.06 $\pm$ 0.32 42.01 $\pm$ 1.12 59.03 $\pm$ 0.39 57.59 $\pm$ 0.76 58.31 $\pm$ 0.46 56.96 $\pm$ 0.53 61.33 $\pm$ 0.39 61.35 $\pm$ 0.41
AUC 75.86 $\pm$ 0.66 72.22 $\pm$ 0.18 74.65 $\pm$ 0.32 64.11 $\pm$ 0.40 76.17 $\pm$ 0.17 72.59 $\pm$ 1.07 72.59 $\pm$ 0.61 74.81 $\pm$ 0.37 78.22 $\pm$ 0.15 77.85 $\pm$ 0.12
25-PHE F1 27.77 $\pm$ 0.91 25.53 $\pm$ 0.11 26.15 $\pm$ 0.12 26.52 $\pm$ 0.29 25.38 $\pm$ 0.52 12.45 $\pm$ 0.54 12.25 $\pm$ 0.55 24.61 $\pm$ 0.71 28.67 $\pm$ 0.33 28.54 $\pm$ 0.25
AUC 63.90 $\pm$ 1.31 62.57 $\pm$ 0.23 70.33 $\pm$ 0.12 56.28 $\pm$ 0.16 72.50 $\pm$ 0.67 55.48 $\pm$ 0.39 58.52 $\pm$ 0.27 71.57 $\pm$ 0.47 74.56 $\pm$ 0.19 74.42 $\pm$ 0.34
Experiment Setting II: In Appendix Table 9 and Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality. In this setting, we empirically find that those methods with missing modality imputation usually shows higher tenacity in high proportion of missingness such as FlexMoE and our propose method. Although ShaSpec has also impute the missing modality by shared information, but it is weak on the backbone model and struggling to optimize with multi-task objectives. Our proposed method, on the other hand, shows superior performance across all setting in CMU-MOSEI and CMU-MOSI for all metrics.
Experiment Setting III: In Appendix Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality. We find that single modality shows very similar performance during inference phrase in our proposed method and Flex-MoE across AUC and F1, indicating the effectiveness of missing modality imputation. In particular, our proposed also shows the highest resistance to missing modality across different combination and remains competitive performance no matter what modality is missing. In addition, the modality combination Audio-Video is the worse modality combination for all implemented models. The potential reason may be that the data interaction between those two modalities are weak or even opposite. We show some visualization evidence in Figure 4 that the embedding of those two modalities are mostly symmetry.
4.3 Why Does Confidence-Guided Gating Improve Performance?
We implemented experiment on switching different gating mechanisms in our propose method as shown in Table 2 and we also plot the expert selection heatmap in Appendix F. Softmax gating mechanism has the worse expert collapse, leading to the poorer performance than the balance softmax gating. However, such load balance loss is hard to learn proper routing scores since the optimization is opposite as stated in Section 2, leading to suboptimal solution in this ablation. Softmax w/ $\mathcal{L}_{load}$ variant shows "Sharp Sinusoidal Wave" expert selection pattern as shown in Figure 3(a), indicating the selection is ambiguous and difficulty on convergence. We also found that considering the uniform weight for all experts can achieve similar performance as sole softmax and generate similar sinusoidal pattern as softmax w/ $\mathcal{L}_{load}$, shown in Figure 9.
(a) Expert selection of softmax w/ $\mathcal{L}_{load}$
(b) Expert selection of ConfNet
Figure 3: Expert selection: X-axis represents epoch and Y-axis represents expert ID. The darker the color, the more selection will be for a particular expert. See detailed and more plots in Appendix F
Table 2: Ablation I: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
Task Metric Mean Softmax Softmax w/ $\mathcal{L}_{load}$ Gaussian Laplacian ConfNet
MIMIC-III F1 48.34 $\pm$ 0.22 48.59 $\pm$ 1.97 51.67 $\pm$ 0.62 46.50 $\pm$ 1.62 50.10 $\pm$ 1.17 53.44 $\pm$ 0.27
AUC 85.13 $\pm$ 0.16 85.91 $\pm$ 0.84 85.97 $\pm$ 0.52 86.70 $\pm$ 0.33 85.47 $\pm$ 1.05 87.05 $\pm$ 0.58
CMU-MOSI F1 42.29 $\pm$ 0.87 41.47 $\pm$ 0.94 42.43 $\pm$ 1.34 42.57 $\pm$ 1.81 43.15 $\pm$ 1.53 44.34 $\pm$ 0.77
AUC 68.01 $\pm$ 1.68 68.77 $\pm$ 1.66 67.40 $\pm$ 0.54 67.74 $\pm$ 0.45 67.60 $\pm$ 1.54 70.41 $\pm$ 1.31
In addition, Laplacian and Gaussian gating function can achieve comparable to softmax w/ $\mathcal{L}_{load}$ gating mechanism without helps of $\mathcal{L}_{load}$, providing further evidence to support the advantage of balance expert selection. Our proposed ConfNet without helps of $\mathcal{L}_{load}$ is the best and significantly outperform other gating function in CMU-MOSI and MIMIC-III. As illustrated in Figure 6 and supported by the quantitative results in Table 2, ConfNet not only preserves expert specialization but also achieves better load balancing than Laplacian, albeit slightly less balanced than softmax with $\mathcal{L}_{\text{load}}$ and Gaussian gating as shown in Appendix F. This comparison provides two insights: First, optimization of $\mathcal{L}_{load}$ is simply against to optimization of softmax gating score but does not help to identify the best expert, reflecting the "Sharp Sinusoidal Wave" pattern in expert selection and ambiguous expert selection. Second, learning specialization of expert can also benefit to downstream performance. Notably, both ConfNet and Laplacian gates demonstrate a common pattern: they maintain the specialization of dominant experts while redistributing some load to less active ones, which highlights that strict load balancing may suppress expert specialization and degrade performance, whereas maintaining a balance between specialization and load distribution leads to superior results.
4.4 How ConfNet and Modality Imputation integrate together?
Table 3: Ablation II: Module dropout. CMU-MOSI adopt 50% missing in Experiment II
Variants MIMIC-III CMU-MOSI
F1 AUC F1 AUC
w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
w/o impute 49.59 $\pm$ 0.29 86.31 $\pm$ 0.12 42.46 $\pm$ 2.29 68.44 $\pm$ 1.82
w/o post-impute 49.86 $\pm$ 1.89 86.53 $\pm$ 0.10 43.14 $\pm$ 1.32 69.13 $\pm$ 1.45
w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 41.98 $\pm$ 0.74 68.48 $\pm$ 1.90
ConfSMoE-T 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 44.34 $\pm$ 0.77 70.41 $\pm$ 1.13
We implement the experiment in Table 3 by removing ConfNet and the two-stage imputation module gradually. This experiment demonstrates that both novel modules, ConfNet and the two-stage imputation, are effective on their own, while their integration provides strong robustness against the missing modality challenge. Specifically, ConfNet captures informative features from the imputed modality, and the adaptive two-stage imputation operation reconstructs high-quality features, thereby complementing each other to achieve robust performance. Specifically, we dropped the entire imputation module and keep the missing modality as zeros (w/o impute), we can see that F1 and AUC dropped off by 3.85 and 0.74 on MIMIC-III, 1.88 and 1.97 on CMU-MOSI. This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design. We further remove the post-imputation (w/o post-impute) and observe that ConfNet can still take advantage of modality-specific feature for their downstream task, showing robustness against missing modality.
4.5 Complexity Analysis
In Table 4, we control the number of parameters for both baseline and proposed methods to compare model complexity and performance. Although ShaSpec and TF achieve lower FLOPs for the same parameter amount, their performance remains less competitive than most Transformer- and MoE-based models, except FuseMoE. Among MoE-based models, ConfSMoE-T delivers the best performance while maintaining similar MFLOPs and parameter counts to the second-best baseline. Under comparable parameter settings, LIMoE and mmFormer fail to keep FLOPs as low as ConfSMoE-T and yield suboptimal results. Moreover, LIMoE is particularly inefficient, as its MFLOPs are substantially higher than those of other models with similar parameter sizes. Overall, ConfSMoE-T achieves the best balance between efficiency and performance among MoE-based baselines.
Table 4: Complexity analysis I: Small Scale
Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE-T
MFLOPs 21,73 4.20 1245.03 237.39 143.93 40.32 39,62
#Params 2,177,406 2,101,571 2,130,227 2,378,371 20,971,020 2,184,963 2,184,730
F1 score 39.12 $\pm$ 1.21 35.81 $\pm$ 0.85 41.11 $\pm$ 0.90 41.14 $\pm$ 0.31 36.52 $\pm$ 0.84 42.12 $\pm$ 1.18 44.63 $\pm$ 1.65
5 Related Works
While in many works considers that the interaction between multimodal data can bring comprehensive understanding for downstream task. Initial multimodal fusion approach usually incoperates neural kernel fusionBucak et al. (2013); Poria et al. (2015), early/late network fusion Xu et al. (2023); Baltrušaitis et al. (2018); Guo et al. (2019). Approaches such as Tensor Fusion NetworkZadeh et al. (2017), Multimodal TransformerTsai et al. (2019) and Multimodal Adaption GateRahman et al. (2020), LlMoEMustafa et al. (2022) all highlights the multimodal interactions are beneficial to downstream task. In medical domain, MedFuse Hayat et al. (2022) incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network. MISTS Zhang et al. (2023) incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality. Yang and Wu (2021) fused multimodal data in a attention gate and compute a replacement vector for modality fusion. However, these approaches are limited to full modality only and difficult to extend to increasing number of modality since the complexity can be significantly increased by the number of modality. Missing modality is realistic and practical problem in real life application such as medical diagnose. SMIL Ma et al. (2021) proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion. Du et al. (2018); Shang et al. (2017) propose generative adversarial networks to impute the missing modality by semi-supervised learning. ShaSpec Wang et al. (2023) and DrFuse Yao et al. (2024) designed multiple loss function to learn share and specific representation of different modality. However, the reconstruction network from SMIL and ShaSpec requires multiple additional auxiliary loss function, which occupies the great amount of gradient in optimization process instead of task specific loss. Sun et al. (2024) In addition, recent studies like FuseMoE Han et al. (2024) respect the missingness of modalities and wish to assign each combination of multimodal data to different experts. Flex-MoEYun et al. (2024) further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.
6 Conclusion
In this work, a novel ConfSMoE model, a confidence-guided sparse mixture-of-experts framework, is proposed to robustly handle multimodal learning with missing modalities. By theoretically analyzing the root cause of expert collapse and empirically validating our findings, we propose two key innovations: a two-stage imputation strategy that preserves modality-specific structure, and a novel confidence-driven gating mechanism that decouples expert selection from softmax-induced sharpness. These components jointly improve both expert specialization and routing stability without requiring entropy-based balancing losses.
Appendix A Notation
Table 5: Notation
Variable Definition
$E_{i}$ i-th Expert Function
E Expert pool
$\textbf{E}_{K}$ Set of Top-K expert
$\mathbf{U}$ ConfNet pool
$\mathbf{h}$ Embedding of Arbitrary Data Input
$K$ Number of K selection
$T$ Number of T selection in Post-Imputation
NN Total number of Expert
$n$ Number of instances selected in Pre-imputation
$G(\cdot)$ Gating Function
$\mathbf{g}$ Gating Score Vector
$\mathcal{H}(\cdot)$ Information Entropy
$p_{t}$ Confidence of Ground Truth
$M_{i,n}$ Modality Pool with i instance and n modality
$t_{s,e}$ Tokens at sequential position $s$ and assigned expert with index $e$
$B$ Sparsity Hyperparameter
$|M|$ Number of Modality
$\mathbf{diag(\cdot)}$ Diagonal Matrix
$\delta_{ij}$ Kronecker delta
$u_{i}$ Softmax input at i-th entry
$c_{i}$ Confidence score for Expert $E_{i}$
$g_{i}$ Gating score for Expert $E_{i}$
To facilitate understanding of the model formulation and derivations, Table 5 outlines the notations used throughout this paper. These include variables associated with expert routing, gating mechanisms, modality assignments, and sparsity control, which are central to our proposed ConfSMoE framework.
Appendix B Derivation of Jacobian for MoE and Load Balance Loss
B.1 Jacobian for Softmax
This section provides detailed mathematical derivations of the Jacobian matrix used in the gating function of ConfSMoE and its relation to the load balance loss. We first derive the Jacobian of the softmax operation applied to the gating logits, which is essential for analyzing expert selection dynamics. We then extend this analysis to derive gradients of the entropy-based load balance loss. These derivations provide theoretical insights into how sparse expert routing and balanced expert utilization can be jointly optimized.
We first derive the Jacobian of the softmax function as used in the MoE formulation Eq. 1.
$\textbf{g}=softmax(u)=\frac{e^{u_{i}}}{\sum^{N}_{j=1}e^{u_{j}}}$ (9)
where $u_{i}$ is the i-th input value of gate logits.
$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} \frac{e^{u_{i}}(\sum^{N}_{j=1}e^{u_{j}}-e^{u_{i}})}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i=j\\ -\frac{e^{u_{i}}e^{u_{j}}}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i\neq j\end{aligned}\right$. (10)
$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} g_{i}(1-g_{j})&\quad\text{if }i=j\\ -g_{i}g_{j}&\quad\text{if }i\neq j\end{aligned}\right$. (11)
Where $g_{i}$ represent $\frac{e^{u_{i}}}{\sum^{N}_{j=1}e^{u_{j}}}$. This can be further compactly represent as:
$\frac{\partial\textbf{g}}{\partial u}=g_{i}(\delta_{ij}-g_{j})$ (12)
Where $\delta_{ij}=\begin{cases}1&\text{if }i=j\\ 0&\text{if }i\neq j\end{cases}$ is Kronecker delta and the matrix form is:
$\textbf{J}_{softmax}=\frac{\partial\textbf{g}}{\partial u}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top}$ (13)
B.2 Jacobian for Load Balance Loss
Assumption 1
For any load balance loss $\mathcal{L}_{\text{load}}$ applied to a Softmax-based MoE, the effect of $\mathcal{L}_{\text{load}}$ is to encourage a flat expert selection distribution, which is equivalent to maximizing the entropy of the gating scores.
With Assumption 1, We have load balance loss $\mathcal{L}_{load}=\frac{1}{\mathcal{H}(\textbf{g})}$, where $\mathcal{H}(\textbf{g})=-\sum^{N}_{i=1}g_{i}log(g_{i})$ is the entropy of the gating distribution.
By applying the chain rule, we obtain:
$\frac{\partial\mathcal{L}_{load}}{\partial\textbf{u}}=-\frac{1}{\mathcal{H}(\textbf{u})^{2}}\frac{\partial\mathcal{H}}{\partial\textbf{u}}=-\frac{1}{\mathcal{H}(\textbf{u})^{2}}\frac{\partial\mathcal{H}}{\partial\textbf{g}}\frac{\partial\textbf{g}}{\partial\textbf{u}}$ (14)
The intermediate gradients are given by:
$\frac{\partial\mathcal{H}}{\partial\textbf{g}}=-log(\textbf{g})-\textbf{g}\frac{1}{\textbf{g}}=-log(\textbf{g})-1,\quad\frac{\partial\textbf{g}}{\partial\textbf{u}}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top}$ (15)
Therefore, he final Jacobian of the load balance loss $J_{load}$ is:
$\mathbf{J}_{load}=\frac{\partial\mathcal{L}_{load}}{\partial\textbf{g}}=[\frac{1}{\mathcal{H}(\textbf{g})^{2}}(log\textbf{g}+1)^{\top}]\cdot(\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top})$ (16)
Appendix C Experiment Details
C.1 Dataset
MIMIC-III Johnson et al. [2016] is a public dataset contains irregular time series and clinical notes modality. We follow the preprocessing scripts from MMIMI-III benchmark Harutyunyan et al. [2019] and Zhang et al. [2023] to form multimodal patient instance. However, Zhang et al. [2023] filter the instances with missing modality and we modified the scripts to preserve with instances with missing modality. MIMIC-IV Johnson et al. [2023] is a public dataset contains irregular time series, clinical notes, ECG signal and chest image. We follow the preprocessing scripts from MedFuse Hayat et al. [2022] to preprocess irregular time series and further preprocess clinical notes, ECG signal and chest image by ourself. We follow the dataset split in MedFuse Hayat et al. [2022] and form multimodal patient instance. CMU-MOSI Zadeh et al. [2016] is a multimodal sentiment analysis dataset comprising audio, video, and text modalities. CMU-MOSEI Bagher Zadeh et al. [2018] is an extension of the CMU-MOSI dataset, utilizing the same modalities—audio, text, and video from YouTube recordings. We follow preprocessing script of CMU-MOSI and CMU-MOSEI from MMMLMustafa et al. [2022].
C.2 Baselines
SMILMa et al. [2021] impute missing modality with Bayesian meta-learning. ShaSpecWang et al. [2023] leverages multiple loss function to learn the modality-share and modality-specific information and impute the missing modality with modality-shared information. TFZadeh et al. [2017] extract the modality interaction by multimodal sub-networks and tensor fusion layer. mmFormerTsai et al. [2019] is a typical multimodal transformer with multihead attention mechanism. LlMoEMustafa et al. [2022] achieves the multimodal learning with contrastive learning. FuseMoEHan et al. [2024] respect the missing modality and jointly learn single and missing modality combination data input and FuseMoE implement Laplacian gating mechanism for modality fusoin and expert selection. In experiment, we denote FuseMoE-S as softmax gating and FuseMoE-L as Laplacian gating. FlexMoEYun et al. [2024] assigns the combination of modality to different experts and impute the missing modality by modality bank. For our approach, we use ConfSMoE-T denotes Token-Level confidence fusion and ConfSMoE-E denotes Expert-level confidence fusion.
C.3 Dataset and Task Description
Table 6: Dataset statistics
Dataset Task #Samples Modality #Classes Is Imbalanced? Missing Ratio
MIMIC-III 48-IHM 6621 Time Series, Text 2 Yes 24.21%
LOS 6621 Time Series, Text 2 Yes 24.21%
25-PHE 12,278 Time Series, Text 25 Yes 24.25%
MIMIC-IV 48-IHM 22,733 Image, ECG, Time Series, Text 2 Yes 47.10%
LOS 22,733 Image, ECG, Time Series, Text 2 Yes 47.10%
25-PHE 47,027 Image, ECG, Time Series, Text 25 Yes 50.32%
CMU-MOSI Sentiment Analysis 1870 Audio, Video, Text 3 Yes 0% - 50%
CMU-MOSEI Sentiment Analysis 20,985 Audio, Video, Text 3 Yes 0% - 50%
48-IHM is a binary classification task, where the time series is truncated to 48 hours readings and predict whether the patient in-hospital mortality in ICU.
LOS is also binary classification task. We formulate task similar to 47-IHM and those patients who spent at least 48 hours in the ICU are filtered to predict the remaining Length Of Stay.
25-PHE is a multilabel classification problem. We attempt to predict one 25 acute care conditions presented in a givem ICU stay record. The definition of those 25 acute care conditions are defined in MIMIC benchmark Johnson et al. [2016]
CMU-MOSI and CMU-MOSEI are multimodal sentiment analysis benchmarks consisting of video clips annotated with sentiment scores. Each sample includes aligned audio, video, and text modalities. To simulate modality missingness, we introduce controlled random dropout during training.
All datasets are publicly available and widely adopted in multimodal research. Those datasets can be found and downloaded from the following links:
  • CMU-MOSI and CMU-MOSEI download link: https://github.com/zehuiwu/MMML
  • MIMIC-III download link: https://physionet.org/content/mimiciii/1.4/ and benchmark: https://github.com/YerevaNN/mimic3-benchmarks
  • MIMIC-IV download link: https://physionet.org/content/mimiciv/2.2/ and benchmark: https://github.com/nyuad-cai/MedFuse
C.4 Hyperparameter and Hardware Setting
All experiments are conducted in a remote server with 4 $\times$ RTX4090 24 GB with 256 GB RAM and 96-cores CPU.
Table 7: Hyperparameter setting
Parameter Value
#Expert 8 for MIMIC-III and MIMIC-IV, 4 for CMU-MOSI and CMU-MOSEI
#Top-K Expert 2
Learning Rate 3e-4
Hidden Size 128
#MoE layers 1
Training Epoch 50
Weight for $\mathcal{L}_{conf}$ 1
#Top-T for Pre-imputation 10
Random Seed 2023, 2024, 2025
Dropout 0.1
Appendix D Limitation
This work may be limited to expert discontinuity as it is build upon on sparse MoE backbone. However, the trade-off between sparse MoE and continuous / fully differentiable MoE, more computational cost vs continuity of experts, is still underexplored.
Appendix E Additional Results
These experiments extend the main results presented in Section Experiment and highlight the scalability of our method to real-world multimodal challenges.
Table 8: Experiment setting I: Main results on MIMIC-III
Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
48-IHM F1 48.56 $\pm$ 0.35 34.36 $\pm$ 0.73 43.80 $\pm$ 0.67 23.09 $\pm$ 1.13 47.62 $\pm$ 0.21 48.62 $\pm$ 0.27 51.63 $\pm$ 0.39 47.81 $\pm$ 0.31 53.44 $\pm$ 0.27 50.00 $\pm$ 2.17
AUC 84.00 $\pm$ 1.10 79.86 $\pm$ 0.53 81.27 $\pm$ 0.76 78.47 $\pm$ 0.65 83.30 $\pm$ 0.64 85.81 $\pm$ 0.13 86.22 $\pm$ 0.15 85.10 $\pm$ 0.61 87.05 $\pm$ 0.58 87.14 $\pm$ 0.21
LOS F1 62.01 $\pm$ 0.54 57.27 $\pm$ 0.36 61.10 $\pm$ 0.65 56.23 $\pm$ 0.79 62.75 $\pm$ 0.10 62.26 $\pm$ 0.36 62.27 $\pm$ 1.61 64.18 $\pm$ 0.34 66.27 $\pm$ 0.40 66.39 $\pm$ 0.23
[…]
25-PHE F1 33.01 $\pm$ 0.30 19.52 $\pm$ 0.39 35.10 $\pm$ 0.38 30.38 $\pm$ 0.38 36.22 $\pm$ 0.39 33.48 $\pm$ 0.22 34.07 $\pm$ 0.51 35.31 $\pm$ 0.57 37.40 $\pm$ 0.13 40.13 $\pm$ 0.40
AUC 75.17 $\pm$ 1.84 64.61 $\pm$ 0.25 71.79 $\pm$ 0.26 66.56 $\pm$ 0.17 77.93 $\pm$ 0.44 77.22 $\pm$ 0.29 77.19 $\pm$ 0.33 77.37 $\pm$ 0.10 78.93 $\pm$ 0.38 78.28 $\pm$ 0.73
Table 9: Experiment setting II: Main results on CMU-MOSI
Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
0% F1 41.78 $\pm$ 0.64 43.28 $\pm$ 0.36 50.22 $\pm$ 0.20 50.91 $\pm$ 0.19 47.69 $\pm$ 0.39 48.15 $\pm$ 0.57 51.50 $\pm$ 0.55 51.83 $\pm$ 0.27 51.70 $\pm$ 0.29
AUC 65.08 $\pm$ 0.39 64.48 $\pm$ 0.51 74.62 $\pm$ 0.49 78.12 $\pm$ 0.31 75.26 $\pm$ 0.20 75.96 $\pm$ 0.29 79.38 $\pm$ 0.23 80.56 $\pm$ 0.10 80.62 $\pm$ 0.07
10% F1 43.40 $\pm$ 0.74 40.14 $\pm$ 0.13 48.66 $\pm$ 1.20 49.91 $\pm$ 0.33 46.81 $\pm$ 0.45 48.81 $\pm$ 0.42 49.92 $\pm$ 0.28 49.99 $\pm$ 0.85 50.89 $\pm$ 0.07
AUC 65.45 $\pm$ 0.56 61.54 $\pm$ 0.66 77.21 $\pm$ 0.76 77.12 $\pm$ 0.14 70.25 $\pm$ 0.93 72.36 $\pm$ 0.50 77.07 $\pm$ 0.21 78.06 $\pm$ 0.64 78.03 $\pm$ 0.21
20% F1 39.15 $\pm$ 0.13 42.64 $\pm$ 0.30 48.51 $\pm$ 1.15 48.84 $\pm$ 0.61 47.92 $\pm$ 0.58 45.00 $\pm$ 0.44 49.16 $\pm$ 0.24 50.58 $\pm$ 0.73 50.17 $\pm$ 0.31
AUC 60.87 $\pm$ 0.29 65.52 $\pm$ 0.98 74.16 $\pm$ 0.71 76.64 $\pm$ 0.23 71.05 $\pm$ 0.36 72.15 $\pm$ 0.37 73.80 $\pm$ 1.43 77.17 $\pm$ 0.55 77.18 $\pm$ 0.11
30% F1 39.54 $\pm$ 0.17 40.42 $\pm$ 0.75 46.66 $\pm$ 1.24 46.13 $\pm$ 0.55 43.78 $\pm$ 0.51 44.48 $\pm$ 0.63 45.49 $\pm$ 0.44 47.57 $\pm$ 0.10 47.52 $\pm$ 0.77
AUC 60.80 $\pm$ 0.44 60.57 $\pm$ 0.99 71.25 $\pm$ 0.57 74.33 $\pm$ 0.49 70.38 $\pm$ 0.56 72.25 $\pm$ 0.48 72.54 $\pm$ 0.92 75.59 $\pm$ 0.18 75.42 $\pm$ 0.49
40% F1 35.32 $\pm$ 0.32 38.69 $\pm$ 0.42 42.22 $\pm$ 0.44 43.31 $\pm$ 0.38 41.88 $\pm$ 0.42 43.20 $\pm$ 0.20 43.13 $\pm$ 0.61 45.12 $\pm$ 0.27 44.18 $\pm$ 0.10
AUC 59.58 $\pm$ 0.13 61.62 $\pm$ 0.07 68.44 $\pm$ 0.14 67.21 $\pm$ 0.43 64.61 $\pm$ 0.63 66.14 $\pm$ 0.11 68.12 $\pm$ 0.34 71.65 $\pm$ 0.47 69.55 $\pm$ 0.49
50% F1 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.14 $\pm$ 0.31 41.54 $\pm$ 0.77 38.55 $\pm$ 0.34 40.90 $\pm$ 0.49 41.95 $\pm$ 0.48 44.34 $\pm$ 0.77 42.48 $\pm$ 0.12
AUC 55.60 $\pm$ 0.12 60.49 $\pm$ 0.80 61.87 $\pm$ 0.67 67.43 $\pm$ 0.34 62.28 $\pm$ 0.28 65.73 $\pm$ 0.30 66.47 $\pm$ 0.49 70.41 $\pm$ 1.31 69.04 $\pm$ 0.32
Table 10: Experiment setting II: Main results on CMU-MOSEI
Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
0% F1 47.48 $\pm$ 0.27 38.89 $\pm$ 0.54 56.75 $\pm$ 0.47 59.13 $\pm$ 0.52 57.04 $\pm$ 0.15 58.71 $\pm$ 0.23 60.62 $\pm$ 0.96 61.31 $\pm$ 0.14 62.35 $\pm$ 0.12
AUC 67.93 $\pm$ 0.56 64.26 $\pm$ 0.61 76.42 $\pm$ 0.24 80.29 $\pm$ 0.31 78.54 $\pm$ 0.42 78.13 $\pm$ 0.34 81.34 $\pm$ 0.20 82.87 $\pm$ 0.10 82.75 $\pm$ 0.44
10% F1 47.41 $\pm$ 0.18 34.94 $\pm$ 1.10 55.37 $\pm$ 0.47 57.46 $\pm$ 1.10 55.90 $\pm$ 0.55 56.68 $\pm$ 0.31 58.66 $\pm$ 1.10 59.81 $\pm$ 0.34 60.22 $\pm$ 0.31
AUC 69.26 $\pm$ 0.22 61.05 $\pm$ 0.34 75.36 $\pm$ 0.53 79.38 $\pm$ 0.94 76.38 $\pm$ 0.39 76.44 $\pm$ 0.12 79.48 $\pm$ 0.23 81.00 $\pm$ 0.10 81.05 $\pm$ 0.17
20% F1 46.54 $\pm$ 1.04 35.37 $\pm$ 0.94 52.91 $\pm$ 0.75 55.35 $\pm$ 0.88 54.44 $\pm$ 0.27 54.60 $\pm$ 0.19 55.96 $\pm$ 1.67 58.06 $\pm$ 0.21 57.66 $\pm$ 0.20
AUC 68.05 $\pm$ 0.56 59.54 $\pm$ 0.22 73.02 $\pm$ 0.65 77.29 $\pm$ 0.47 74.32 $\pm$ 0.16 74.92 $\pm$ 0.64 78.05 $\pm$ 0.10 78.85 $\pm$ 0.28 78.83 $\pm$ 0.22
30% F1 45.05 $\pm$ 0.90 35.36 $\pm$ 0.84 49.71 $\pm$ 0.30 53.93 $\pm$ 1.13 52.27 $\pm$ 0.35 49.92 $\pm$ 0.36 52.00 $\pm$ 0.47 54.16 $\pm$ 0.49 54.83 $\pm$ 0.15
AUC 66.89 $\pm$ 0.31 60.23 $\pm$ 0.31 71.28 $\pm$ 0.10 76.50 $\pm$ 0.31 72.22 $\pm$ 0.52 72.52 $\pm$ 0.29 76.14 $\pm$ 0.38 76.91 $\pm$ 0.41 77.10 $\pm$ 0.30
40% F1 43.49 $\pm$ 0.95 32.23 $\pm$ 0.61 48.06 $\pm$ 0.54 50.95 $\pm$ 0.99 49.32 $\pm$ 0.17 49.36 $\pm$ 0.60 49.29 $\pm$ 0.28 50.72 $\pm$ 0.37 50.66 $\pm$ 0.18
AUC 65.32 $\pm$ 0.59 60.09 $\pm$ 0.35 69.31 $\pm$ 0.10 73.24 $\pm$ 0.76 69.85 $\pm$ 0.11 70.05 $\pm$ 0.46 72.08 $\pm$ 1.27 73.63 $\pm$ 0.48 73.88 $\pm$ 0.25
50% F1 42.75 $\pm$ 1.22 33.01 $\pm$ 0.51 47.39 $\pm$ 0.16 46.69 $\pm$ 0.63 43.08 $\pm$ 0.47 44.44 $\pm$ 0.36 46.19 $\pm$ 0.34 47.31 $\pm$ 0.20 48.13 $\pm$ 0.33
AUC 62.32 $\pm$ 0.42 58.20 $\pm$ 0.17 66.61 $\pm$ 0.22 70.90 $\pm$ 0.82 67.63 $\pm$ 0.55 67.44 $\pm$ 0.16 70.15 $\pm$ 1.10 71.74 $\pm$ 0.14 70.48 $\pm$ 0.11
Table 11: Experiment setting III: Main results on CMU-MOSI
Test Modality Dataset: CMU-MOSI / Evaluation Metric: AUC
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
62.18 $\pm$ 0.88 72.72 $\pm$ 0.08 59.37 $\pm$ 1.12 74.63 $\pm$ 0.13 55.35 $\pm$ 1.11 56.44 $\pm$ 1.48 74.59 $\pm$ 1.50 75.57 $\pm$ 0.55 75.56 $\pm$ 0.33
55.79 $\pm$ 2.53 52.21 $\pm$ 0.31 57.87 $\pm$ 0.32 56.84 $\pm$ 0.34 56.24 $\pm$ 1.96 55.11 $\pm$ 0.77 52.28 $\pm$ 1.23 58.42 $\pm$ 0.14 58.66 $\pm$ 0.20
65.22 $\pm$ 1.01 71.96 $\pm$ 0.47 57.89 $\pm$ 0.65 75.59 $\pm$ 0.39 54.36 $\pm$ 1.67 56.10 $\pm$ 1.76 75.55 $\pm$ 1.64 77.02 $\pm$ 0.63 76.96 $\pm$ 0.64
55.10 $\pm$ 1.59 71.01 $\pm$ 0.53 58.57 $\pm$ 0.44 74.73 $\pm$ 0.51 55.94 $\pm$ 1.51 55.66 $\pm$ 1.66 74.35 $\pm$ 1.60 75.07 $\pm$ 0.70 75.11 $\pm$ 0.54
61.60 $\pm$ 4.80 73.77 $\pm$ 0.44 60.28 $\pm$ 1.42 75.73 $\pm$ 0.33 55.45 $\pm$ 0.95 56.58 $\pm$ 1.29 74.42 $\pm$ 1.39 75.58 $\pm$ 0.55 75.50 $\pm$ 0.41
66.85 $\pm$ 3.01 69.04 $\pm$ 0.94 57.19 $\pm$ 0.72 73.29 $\pm$ 0.53 56.04 $\pm$ 1.91 57.85 $\pm$ 0.74 75.34 $\pm$ 1.46 77.02 $\pm$ 0.63 76.47 $\pm$ 0.63
Table 12: Experiment setting III: Main results on CMU-MOSI
Test Modality Dataset: CMU-MOSI / Evaluation Metric: F1
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
41.27 $\pm$ 0.81 48.25 $\pm$ 0.11 34.96 $\pm$ 0.72 44.45 $\pm$ 0.33 30.51 $\pm$ 2.53 29.90 $\pm$ 2.58 50.54 $\pm$ 0.73 54.07 $\pm$ 0.29 54.04 $\pm$ 0.36
36.99 $\pm$ 2.03 35.79 $\pm$ 0.55 34.83 $\pm$ 0.54 35.20 $\pm$ 0.64 24.29 $\pm$ 4.23 27.89 $\pm$ 6.23 30.02 $\pm$ 7.82 37.40 $\pm$ 0.16 38.17 $\pm$ 0.13
43.72 $\pm$ 1.64 45.53 $\pm$ 0.33 37.25 $\pm$ 0.53 47.79 $\pm$ 0.46 20.68 $\pm$ 2.14 29.88 $\pm$ 4.55 50.41 $\pm$ 1.55 52.86 $\pm$ 0.75 51.13 $\pm$ 0.15
42.98 $\pm$ 0.21 46.49 $\pm$ 0.91 38.11 $\pm$ 0.35 44.86 $\pm$ 0.94 25.67 $\pm$ 2.82 27.34 $\pm$ 2.94 50.39 $\pm$ 0.87 53.76 $\pm$ 0.43 54.57 $\pm$ 0.35
41.36 $\pm$ 2.27 47.13 $\pm$ 0.54 37.99 $\pm$ 0.40 43.81 $\pm$ 0.42 25.70 $\pm$ 2.82 27.30 $\pm$ 2.95 50.65 $\pm$ 0.90 53.77 $\pm$ 0.45 49.93 $\pm$ 0.06
45.45 $\pm$ 2.73 46.02 $\pm$ 1.21 38.79 $\pm$ 0.56 45.50 $\pm$ 0.47 25.23 $\pm$ 4.59 29.73 $\pm$ 3.51 50.57 $\pm$ 1.35 53.72 $\pm$ 0.94 50.69 $\pm$ 0.47
Table 13: Experiment setting III: Main results on CMU-MOSEI
Test Modality Dataset: CMU-MOSEI / Evaluation Metric: AUC
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
64.47 $\pm$ 0.10 76.60 $\pm$ 0.39 56.98 $\pm$ 0.20 81.23 $\pm$ 0.07 75.81 $\pm$ 0.82 76.62 $\pm$ 0.37 80.63 $\pm$ 0.41 81.45 $\pm$ 0.02 81.52 $\pm$ 0.07
58.41 $\pm$ 0.10 56.72 $\pm$ 0.84 58.41 $\pm$ 0.17 59.64 $\pm$ 0.56 57.32 $\pm$ 0.62 57.22 $\pm$ 0.12 58.28 $\pm$ 1.51 60.33 $\pm$ 0.03 60.00 $\pm$ 0.23
64.46 $\pm$ 0.14 75.91 $\pm$ 0.04 58.64 $\pm$ 0.13 80.00 $\pm$ 0.29 77.24 $\pm$ 0.60 77.31 $\pm$ 0.16 80.00 $\pm$ 0.10 81.26 $\pm$ 0.10 81.23 $\pm$ 0.14
63.52 $\pm$ 0.19 75.82 $\pm$ 0.28 58.71 $\pm$ 0.35 80.69 $\pm$ 0.34 76.84 $\pm$ 0.14 76.27 $\pm$ 0.70 81.06 $\pm$ 0.20 81.60 $\pm$ 0.13 81.53 $\pm$ 0.22
64.40 $\pm$ 0.24 75.83 $\pm$ 0.28 58.89 $\pm$ 0.40 80.96 $\pm$ 0.26 76.85 $\pm$ 0.14 76.51 $\pm$ 0.50 80.48 $\pm$ 0.21 81.49 $\pm$ 0.13 81.65 $\pm$ 0.17
65.16 $\pm$ 0.04 75.85 $\pm$ 0.09 58.32 $\pm$ 0.04 80.19 $\pm$ 0.11 77.13 $\pm$ 0.75 77.23 $\pm$ 0.24 80.21 $\pm$ 0.11 81.25 $\pm$ 0.10 81.23 $\pm$ 0.14
Table 14: Experiment setting III: Main results on CMU-MOSEI
Test Modality Dataset: CMU-MOSEI / Evaluation Metric: F1
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
40.75 $\pm$ 0.14 56.02 $\pm$ 0.56 31.38 $\pm$ 0.61 58.97 $\pm$ 0.53 55.85 $\pm$ 0.69 53.59 $\pm$ 1.07 60.10 $\pm$ 0.76 61.56 $\pm$ 0.22 61.78 $\pm$ 0.58
38.15 $\pm$ 0.92 37.12 $\pm$ 0.66 31.73 $\pm$ 0.16 38.26 $\pm$ 2.40 29.40 $\pm$ 1.67 33.29 $\pm$ 2.89 36.28 $\pm$ 3.51 39.76 $\pm$ 0.23 39.36 $\pm$ 1.45
40.82 $\pm$ 0.12 56.06 $\pm$ 0.29 31.86 $\pm$ 0.23 58.58 $\pm$ 0.31 57.02 $\pm$ 1.60 53.38 $\pm$ 0.65 58.70 $\pm$ 2.10 61.37 $\pm$ 0.28 61.15 $\pm$ 0.39
40.37 $\pm$ 0.46 55.70 $\pm$ 0.35 31.59 $\pm$ 0.23 61.20 $\pm$ 0.46 54.02 $\pm$ 2.88 53.08 $\pm$ 0.40 58.58 $\pm$ 0.71 61.84 $\pm$ 0.70 61.78 $\pm$ 0.58
42.16 $\pm$ 0.22 55.89 $\pm$ 0.37 30.84 $\pm$ 0.26 59.35 $\pm$ 0.12 55.87 $\pm$ 0.68 54.11 $\pm$ 0.49 58.42 $\pm$ 0.27 61.56 $\pm$ 0.61 61.94 $\pm$ 0.37
42.45 $\pm$ 0.85 56.06 $\pm$ 0.29 30.71 $\pm$ 0.71 60.00 $\pm$ 0.85 56.64 $\pm$ 1.65 52.92 $\pm$ 1.11 59.50 $\pm$ 0.80 61.10 $\pm$ 0.10 61.00 $\pm$ 0.76
Appendix F Visualization of Experiment
In this section, we provide visual evidence to further support our claims regarding expert selection dynamics and modality robustness.
Figure 4 shows the UMAP projection of multi-modal embeddings on CMU-MOSI. It clearly demonstrates that imputed modalities form distributions consistent with their original counterparts, suggesting that our imputation module maintains semantic alignment across modalities.
Figure 4: UMAP Visualization of Multi-modal Embedding Space: CMU-MOSI
Figure 5 illustrates the softmax gating mechanism without regularization, which exhibits severe expert collapse where only a few experts are consistently selected throughout training. In contrast, Figure 6 shows that our proposed ConfNet gating maintains both load balance and expert specialization, effectively mitigating collapse and enabling more stable and diverse expert routing.
Figure 5: Softmax Score-based Expert Selection: Serious Expert Collapse when training goes on
Figure 6: ConfNet Gating Expert Selection: Expert collapse is consistently mitigated throughout training. While certain experts are preferred, less frequently selected experts are still actively utilized. Moreover, expert selection becomes increasingly balanced as training progresses, indicating improved stability and diversity in routing.
Figures 7 to 10 compare various gating mechanisms under similar conditions.
Figure 7: Softmax Gate with Load Balance Loss Expert Selection: An optimization conflict arises, often reflected in a "Sharp Sinusoidal Wave" pattern, experts that were heavily selected in one epoch tend to be suppressed in the next. This oscillatory behavior indicates that the typical load balancing loss does not promote true specialization or expertise among experts. Instead, it merely enforces uniform usage, potentially at the cost of performance, by prioritizing balanced selection over actual expert quality.
Figure 8: Laplacian Gating Expert Selection: Laplacian gating can partially alleviate the load imbalance issue and achieves better expert utilization than softmax. However, as training progresses, selection becomes increasingly skewed, with only a few experts being consistently chosen. This eventually leads to expert collapse, where most experts become inactive and underutilized.
Figure 9: Mean Gating Expert Selection: The "Sharp Sinusoidal Wave" pattern persists under mean gating, where experts heavily selected in one epoch are suppressed in the next. This oscillatory behavior reflects ambiguity in expert selection, stemming from the absence of explicit guidance in the routing mechanism. While mean gating achieves better load balance than Softmax with $\mathcal{L}_{\text{load}}$, it ultimately results in inferior performance, as shown in Table 2.
Figure 10: Gaussian Gating Expert Selection: The "Sharp Sinusoidal Wave" pattern persists under Gaussian gating, where experts heavily selected in one epoch are suppressed in the next. This oscillatory behavior reflects ambiguity in expert selection, stemming from the inappropriate gating score guided from Gaussian. While mean gating achieves better load balance than Softmax with $\mathcal{L}_{\text{load}}$, it ultimately results in inferior performance, as shown in Table 2.
Finally, Figures 11 and 12 report Experiment III results. ConfMoE demonstrates superior resilience across increasing missing rates, with both F1 and AUC metrics degrading gracefully. These results visually confirm the effectiveness of confidence-guided gating and modality imputation under challenging conditions.
Figure 11: Experiment III: CMU-MOSEI Performance Plot
Figure 12: Experiment III: CMU-MOSI Performance Plot
Figure 13: Attention Map from post-imputation: Text as query and Vision, Text as key. The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token. The darker the color, the lower attention weight it is.
Figure 14: Attention Map from post-imputation: Vision as query and Audio, Text as key. The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token. The darker the color, the lower attention weight it is.
Figure 15: Attention Map from post-imputation: Audio as query and Vision, Text as key. The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token. The darker the color, the lower attention weight it is.
From Figure 13 14 15, we can clearly observed that the attention map are distinctly different when the key are set to different modalities. This indicate that the post-imputation module can dynamically learn, which token are needed depending on characteristics of token and experts. For example, the Figure 15 indicates that 25 - 45th tokens of the first expert are more important than others while the text modalities shows that the second experts are more important. Similar observation are also shows in 13 14 as well. Therefore, we can conclude that the post-imputation can capture expertise of different expert and refine the missing modality accordingly.